A 555 mile, 5 hour plane trip was flown at two speeds. For the first part of the trip, the average speed was 105mph. For the remainder of the trip, the average speed was 115mph. How long did the plane fly at each speed?
Answer by lwsshak3(11628)

A 555 mile, 5 hour plane trip was flown at two speeds. For the first part of the trip, the average speed was 105mph. For the remainder of the trip, the average speed was 115mph. How long did the plane fly at each speed?
..
let x= hours of travel time during first part of trip
5-x= hours of travel time during remainder of trip
distance = speed*travel time
105x+115(5-x)=555
105x+575-115x=555
10x=20
x=2
5-x=3
..
Travel time during first (115 mph) part of trip=2 hrs
Travel time during remaining (105 mph) part of trip=3hrs